

********************************************************************
UARS WEEKLY STATUS REPORT - 21 MAY 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM gathered its normal solar, offset, and occultation data without
apparent problems.
***********
PEM
***********
PEM/AXIS was active from May 14 through May 18. During this time,
PEM/AXIS generated deposition profiles from electron precipitation which
were inverted from measured X-rays.
PEM, active under PMON control, was terminated on May 18 due to a
large energy drain on the UARS spacecraft. While under PMON control, the
PEM instrument computed energy deposition from its particle spectrometers.
The FOT and the PEM team are completing close-out documentation on
the HEPS Energy Processor trouble; however progress has been hampered
with the departure of Peter Thomas.
***********
HALOE
***********
HALOE has been successfully operating. Presently, HALOE is off, but is
scheduled to be powered on beginning of day May 29, 1999.
***********
MLS
***********
MLS was reactivated successfully on May 11 and operated until May 19 13:27 UTC
when the FOT needed to turn MLS off due to power constraints. MLS was reactived
on May 20 and is scheduled to operate through May 28.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
2 days CLA (7.0) PROD (L1-L3) (22R-12C-4F=6L)
2 days HAL (4.81) PROD (L1-L3) (2R-2C=0L)
2 day HRD (4.4) PROD (L1-L3)(TEST) (2R-2C=0L)
50 days PEM (3.32) PROD (OA-L2) (753R-718C-1F=34L)
55 days WIN (5.11) PROD (LOG) (1540R-1540C=0L)
62 days WIN (5.11) PROD (L1-L3) (1540R-512-4F=1024L)
March 1999 current day DAAC data will be loaded in May 1999.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 08 May 1999 (Orbit 41830, GMT Day 128) through 14 May
1999 (GMT Day 134).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
ACRIM 130/2230 41873 Off
MLS 131/0006 41874 On
SOLSTICE 131/0015 41874 Day Only
HRDI 131/0336 41876 Day Only
HALOE 133/2246 41918 Off
SOLSTICE 134/0025 41919 Full Time
ACRIM 134/0059 41919 On
WINDII 134/0059 41919 On
PEM 134/1126 41926 AXIS On
The beta angle ranged between the angles of 26.0 degrees to 55.9 degrees
for this report period. The beta angle is now increasing toward a maximum
angle of 77 degrees on 21 May 1999 (DOY 141).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.7 to 1.3 degrees C. Additional battery performance data
for each battery on May 07 1999 and May 14 1999 are provided in Appendix D.
The clock error ranged from 11.1 to -11.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% ACRIM 02
% PEM 02
% MLS 01
% HRDI 01
% HALOE 01
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
% HRDI 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/117 - 2287.496987
Transponder B: 99/118 - 2287.498739
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/119 show the observatory data loss to be
24 hours, 12 minutes, 39.348 seconds (an increase of 2.240 seconds since
the last report period). This is a 0.0361 percent data loss which equals a
99.9639 percent data capture for the mission.
The increase of 2.240 seconds occurred over five (3) days:
7 1.056 seconds on Day 127 due to a data hit,
7 0.672 seconds on Day 132 due to a data hit,
7 0.512 seconds on Day 133 due to a data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The
current work is a continuation of the SCS-21 UARS implementation effort
done last year which culminated in a level 1 demonstration.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
The end-to-end testing that was scheduled to start this week has been
delayed until next week to allow additional time for FDF to create special
Y2K planning aids.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL - No change
MEETINGS
FOT attended the weekly IMOC meetings on 12 May 1999.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (forward-to-reverse) is scheduled for 11
June 1999 (DOY 162).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 14 May 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
41845 129/0139 99-159 OBC Clk Adj fr 0.00 to 20.6 msec/day
41871 130/1942 99-160 OBC Clk Adj fr 20.6 to 0.00 msec/day
41873 130/2230 99-161 ACRIM Power OFF
41874 131/0006 99-162 MLS Power ON
41876 131/0336 99-163 HRDI Day only operations
41903 132/2228 99-164 OBC Clk Adj fr 0.00 to 20.6 msec/day
41918 133/2246 99-165 HALOE Power OFF
41919 134/0059 99-166 ACRIM Power ON
41919 134/0059 99-167 WINDII Power ON
41926 134/1126 99-168 PEM AXIS 1&2 ON
41930 134/1747 99-169 PEM PEM CALS
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 14 May 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIISUTEMP Y-LO WINDII ISU Temp 41881-88
WIISUTEMP Y-LO WINDII ISU Temp continuous
SORLGCTMP Y-LO Logic_Box Temperature continuous
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 14 May 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
41845 129/01:39 (08 May 99) from 00.0 to 20.6 msec/day
41871 130/19:42 (10 May 99) from 20.6 to 00.0 msec/day
41903 132/22:28 (12 May 99) from 00.0 to 20.6 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 14 May 1999
07 May 1999 - GMT Day 127 Beta = 26.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -78.4/-100.8 -3.05/-4.31 +0.0/ +0.0 16.5 0.000 0.0
2 +39.2/-16.8 +6.11/+4.83 +30.4/-15.2 24.8 1.023 22.1
3 +22.4/-50.4 +5.15/+3.87 +30.4/-16.0 24.8 1.024 22.0
14 May 1999 - GMT Day 134 Beta = 55.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -78.4/-112.0 -3.68/-4.94 +0.0/ +0.0 16.5 0.000 0.0
2 +22.4/-28.0 +5.15/+3.87 +24.8/-16.4 23.8 1.013 21.0
3 +22.4/-50.4 +4.51/+2.60 +24.8/-17.2 23.8 1.016 20.8
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 14 May 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-026 HRDI 41875 131/0148 N/A Not in Day Only as Scheduled
99-023 FOT 41929 134/1604 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-024 FOT 41741 122/0302 20809 Late Acquisition
99-026 HRDI 41875 131/0148 N/A Not in Day Only as Scheduled
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-023 FOT see list May 99 14170 Generic Late Acq
SEVNTFW99/20